# Fair Graph Machine Learning under Adversarial Missingness Processes

### v4

@@ 4.2 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are provided in Appendix A.
@@
The optimal $f^{*}_{I}$ generates the worst-case (minimum fairness) imputation so that $\Delta DP_{f_{C},f^{*}_{I}}\geq\Delta DP_{f_{C},f_{I}}$ where, $\Delta DP_{f_{C},f_{I}}=|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$, $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ and, $\hat{s}=f_{I}(\mathcal{G},\mathcal{X}).$
$f_{I}$ generates imputations such that $\Delta DP$ in Sec 5) is maximized for a given classifier $f_{C}$ .
@@
The maximal $\Delta DP$ of BFtS is minimal among all estimators of $s$.
The optimal GNN classifier $f^{*}_{C}$, will achieve demographic parity ($\Delta DP=0$ in Sec 5) for the worst-case imputation (minimum fairness) $\hat{s}$ generated by $f_{I}$.
@@
Let $s^{\prime}=f(G,X)$ be an imputation method independent of models $f_{C}$ and $f_{A}$, then the Jensen Shannon divergence $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\leq JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is related to the convergence of adversarial learning. For FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eq. 4 will be constant. The interplay between the three players makes BFtS more robust to convergence issues because the objective minimizes the upper bound on the JS divergence. This reduces the probability that the divergence vanishes during training (see details in Appendix A).
5 Experiments and Results
@@
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic data. The size of $\mathcal{V}_{S}$ was set to $30\%$ of $|\mathcal{V}|$ unless stated otherwise. The details of the datasets, baselines, hyperparameters, additional experiments with graph attention network, and ablation studies are in Appendix A.
Figure 6: Fairness vs. accuracy for all the methods. The x-axis of each plot shows the F1 score. We plot $1-\Delta DP$ and $1-\Delta EQOP$ on the y-axis of rows 1 and 2, respectively. Therefore, the top right corner of the plot represents a high F1 with low bias. Our approach (BFtS) often achieves better fairness for a similar value of F1 compared to the baselines.
@@
Table 1: AVPR, F1 score, $\%\Delta DP$, and $\%\Delta EQOP$ without any sensitive information of BFtS and RNF. Amongst the baselines, only RNF works without any sensitive information. BFtS outperforms RNF in terms of fairness and accuracy for all baselines.
Evaluation: For utility evaluation, we use average precision (AVPR) and F1 score. For measuring bias we use $\Delta DP$ and $\Delta EQOP$. Details of these metrics are in Appendix A.
@@ 5.1 Results and Analysis
Figure 5 shows the F1 score, $1-\Delta DP$, and $1-\Delta EQOP$ for different methods while varying their hyperparameter. We have generated the simulation graph using a stochastic block model with different assortativity coefficients, i.e., the extent to which links exist within clusters compared with across clusters. Learning missing sensitive values and node labels with a low assortativity coefficient is hard and graphs with assortativity $0.17$ and $0.37$ represent the scenario described in Corollary 1, therefore, the JS divergence tends to be small for independent imputation, which may result in a lack of convergence for the baselines. Figure 5 shows that for assortativity $0.17$ and $0.37$, FairGNN and Debias fail to predict node labels in multiple cases. As we increase the assortativity, all baselines can predict the labels, but BFtS still achieves a better fairness vs. accuracy trade-off in all cases.
We demonstrate the accuracy vs. fairness trade-off for real datasets in Figure 6. For BFtS, we vary $\alpha$ and $\beta$ to generate different values of fairness and accuracy. For other models, we vary their respective regularizers. The top right corner of the plot represents high fairness and accuracy. Our model achieves better fairness and similar accuracy to the best baseline for the Bail, NBA, German, pokec-n and pokec-z. For all datasets, BFtS achieves a better fairness-accuracy tradeoff.
Table 1 shows the accuracy and fairness results for Bail, Credit, German, Pokec-z, Pokek-n and NBA without any sensitive attribute information. Amongst the baselines, only RNF works in this setting. The results show that BFtS outperforms RNF for the Bail, German, and Credit. Our approach also outperforms RNF in terms of fairness on Pokec-z,Pokec-n, and NBA with similar AVPR and F1. As expected, we note that when we have $30\%$ of the sensitive attributes available (Figure 6), we obtain higher fairness than when we have no sensitive attribute information (Table 1).
@@ 6 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of adversarially missing sensitive attributes that produce challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate BFtS using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.
Limitations and societal impact: We note the following as limitations of our work: (1) our experiments do not consider large-scale graph datasets for fairness as they are currently unavailable, and (2) our theoretical analysis shows that BFtS achieves demographic parity at equilibrium but not other fairness notions. In terms of social impact, our approach enables fair graph ML with missing data. However, one can leverage our method to predict private information.
@@ Appendix A Technical Appendices and Supplementary Material
Two of the models (or players) learned by our model are GNNs. Without loss of generality, we will assume that they are message-passing neural networks [25], which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
@@
Fairness Evaluation: In machine learning literature, fairness is often measured with bias. We apply two bias metrics that will be defined based on a sensitive attribute $s$. Let $s=1$ be the sensitive group and $s=0$ be the non-sensitive group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute [36]. We evaluate demographic parity as:
@@
Equality of Opportunity (EQOP) calculates the difference between true positive rates between sensitive and non-sensitive groups [28] as follows:
@@
Bail [34]: Contains 18,876 nodes representing people who committed a criminal offense and were granted bail by US state courts between 1990 and 2009. Defendants are linked based on similarities in their demographics and criminal histories. The objective is to categorize defendants as likely to post bail vs. no bail. No-bail defendants are expected to be more likely to commit violent crimes if released. Race information is considered a sensitive attribute. Features include education, marital status, and profession.
Credit [74]: The dataset has $30,000$ nodes representing people each with $25$ features, including education, total overdue counts, and most recent payment amount. The objective is to forecast if a person will or won’t miss a credit card payment while considering age as a sensitive attribute. Nodes are connected based on payment patterns and spending similarities.
German [30]: The dataset consists of 1000 nodes and 20 features (e.g., credit amount, savings, profession) representing clients at a German bank. Labels are credit decisions for a person (good or bad credit). The sensitive attribute is gender. The goal is to predict who should receive credit.
NBA [14]: Dataset based on a Kaggle dataset about the National Basketball Association (US). It contains 403 players as nodes connected based on their Twitter activity. Attributes represent player statistics and nationality is considered the sensitive attribute. The goal is to predict the players’ salaries (above/under the median).
Pokec [63]: This dataset has two variations: Pokec-z and pokec-n. The pokec dataset is based on a popular social network in Slovakia. The sensitive attribute in pokec-z and pokec-n are two different regions of Slovakia. The goal is to predict the working field of users.
@@
Debias [77]: Debias is an adversarial learning method that trains a classifier and an adversary simultaneously. The adversary is trained using the softmax output of the last layer of the classifier. It does not handle missing sensitive data. Hence, we only use $\mathcal{V}_{S}\in\mathcal{V}$ to calculate the adversarial loss.
FairGNN [14]: FairGNN first uses a GNN to estimate the missing sensitive attributes. It then trains another GNN-based classifier and a DNN adversary. The adversary helps to eliminate bias from the representations learned by the GNN classifier.
RNF [17]: RNF imputes the missing sensitive attributes by first training the model with generalized cross entropy [80]. It eliminates bias from the classification head. RNF uses samples with the same ground-truth label but distinct sensitive attributes and trains the classification head using their neutralized representations.
FairVGNN [69]: FairVGNN learns fair node representations via automatically identifying and masking sensitive-correlated features. It requires complete sensitive information, so we use the same imputation method as FairGNN for the missing values.
FairSIN [73]: FairSIN learns fair representations by adding additional fairness-facilitating features. It also requires complete sensitive information, so we use the same independent imputation method as FairGNN for the missing values. [73]:
@@
Each GNN has two layers and a dropout probability of $0.5$. In practice, some sensitive information is often available, unless otherwise stated. Therefore, we concentrate the majority of our experiments on the assumption that there is a limited amount of sensitive information available, and we generate the imputations using the available sensitive information in Equation 2. The $\mathcal{V}_{S}$ was set to $30\%$ for all experiments unless stated otherwise. $\mathcal{V}_{L}$ is $300,400,1400,100,500,500$, and $800$ for the Simulation, German, Credit, NBA, pokec-z, pokec-n and Bail datasets, respectively. We adjust the values of $\alpha$ and $\beta$ for each dataset using cross-validation based on the F1 score. For all baselines, we choose the respective regularizers using cross-validation based on the same metrics. Every model, except RNF, combines the training of several networks with various convergence rates. We stop training RNF based on early stopping to eliminate overfitting. We train the rest of the models for certain epochs and choose the model with the best AVPR. All the results shown in the following section are the average of $10$ runs. We use GCN as the GNN architecture for all experiments unless mentioned otherwise.
@@
We will consider an adversarial active learning setting, where the goal of the adversary is to select $k$ labels (i.e., observed sensitive values) to minimize the bias $\mathcal{L}_{A}$ of the data imputed by $f_{I}$. More specifically, our proof will use the coverage model for active learning, where an unlabeled data point $x_{i}^{u}$ can be correctly predicted iff it is covered by at least one labeled point $x_{j}^{l}$, so the goal is for the labeled points $\{x_{1}^{l},x_{2}^{l},\ldots x_{k}^{l}\}$ to cover as many unlabeled data points $\{x_{1}^{u},x_{2}^{u},\ldots x_{n-k}^{u}\}$ as possible [75, 57]. Thus, in our setting, the adversary’s goal is to select labels to minimize the coverage of the following sets of nodes: (1) nodes in the sensitive group ($s_{v}=1$) and the negative class ($y=0$) and (2) nodes in the non-sensitive group ($s_{v}=0$) and the positive class ($y=1$). Without loss of generality, we will focus on the second set $(s_{v}=0,y=1)$ because it impacts demographic parity as a bias metric. This can be achieved by minimizing the coverage of such nodes. We provide a reduction from the minimum k-union problem.
Minimum k-union problem (MkU, [13]): Given a set of sets $\mathcal{S}={S_{1},S_{2},\ldots S_{q}}$, where each set $S_{i}\subseteq\mathcal{I}$ and $\mathcal{I}$ is a set of items. The problem consists of selecting $r$ sets $S_{1},S_{2},\ldots S_{r}$ from $\mathcal{S}$ to minimize the coverage $S_{1}\cap S_{2}\cap\ldots S_{r}$.
Under our coverage model, the reduction is straightforward.
Given an instance of MkU, we define a corresponding instance of the Adversarial Missingness for Minimizing the Bias in the Data (AMMBD) problem as follows. We represent each item in $\mathcal{I}$ and each set in $\mathcal{S}$ as a data point $x$ and assign all data points to both the non-sensitive group ($s_{v}=0$) and the positive class ($y=1$). Moreover, we will add one data point $x^{\prime}$ that will belong to the sensitive group ($s_{v}=1$) and the positive class ($y=1$). Therefore, the resulting set of data points is $\{x_{1},x_{2},\ldots x_{n}\}$, where $n=|\mathcal{S}|+|\mathcal{I}|+1$. Moreover, let $x_{i}$ cover $x_{j}$ iff $x_{i}$ represents a set $S_{i}\in\mathcal{S}$, $x_{j}$ represents an item $i_{j}\in\mathcal{I}$, and $i_{j}\in S_{i}$. It follows that minimizing the bias in the AMMBD instance by selecting data points $x_{j}$ to be labeled is equivalent to minimizing the coverage in the corresponding MkU instance.
In terms of approximation, the best-known approximation algorithm for the minimum k-union problem used in our reduction is $O(n^{1/4+\epsilon})$, where $n$ is the number of sets and $\epsilon>0.$, which is mapped to the size of the graph in our problem [13]. This means that at least $O(kn^{1/4})$ nodes would have to be selected to achieve the optimal coverage for $k$ nodes. We note that this is a worse approximation than the classical $O(\log(n))$-approximation for set cover achieved by a greedy algorithm.
@@
$\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}$ $[\log(f_{A}(\mathbf{h}))]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$
$=\sum_{h}p(\mathbf{h}|\hat{s}=1)\log(f_{A}(\mathbf{h}))+\sum_{h}p(\mathbf{h}|\hat{s}=0)\log(1-f_{A}(\mathbf{h}))$
If we fix $\theta_{C}$ and $\theta_{I}$, then the optimal $f_{A}^{*}$ is $\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}.$ By substituting the optimal $f_{A}^{*}$ in Eq. 7 we get:
$\min_{\theta_{C}}\max_{\theta_{I}}$~ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}]+$
@@
By removing the constants, we can further simplify the min-max optimization with optimal $f_{A}^{*}$ to:
$\min_{\theta_{C}}\max_{\theta_{I}}JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$
The optimal $f_{I}^{*}$ thereby maximizes the following objective:
@@
Since $\hat{y}=\sigma(\mathbf{h.w})$, $f^{*}_{I}$ generates imputation so that
$\min_{\theta_{C}}\sup_{\hat{s}}JS(p(\hat{y}|\hat{s}=1);p(\hat{y}|\hat{s}=0))$
The supremum value will make the two probabilities maximally different. Therefore, for the optimal $f^{*}_{I}$ we get,
$\min_{\theta_{C}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Let $\Delta DP_{f_{C},f_{I}}=|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$, then, the optimal $f^{*}_{I}$ generates the worst-case (minimum fairness) imputation so that $\Delta DP_{f_{C}.f^{*}_{I}}\geq\Delta DP_{f_{C},f^{*}_{I}}$

### v5

@@ 4.2 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are provided in Appendix.
@@
BFtS learns an imputation model $f_{imp}$ with the worst-case imputation:
$\theta_{imp}^{*}=\arg\max_{\theta_{imp}}\mathcal{L}_{bias}=\arg\max_{\theta_{imp}}|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$
where $\hat{y}=f_{class}(\mathcal{G},\mathcal{X})$ and, $\hat{s}=f_{imp}(\mathcal{G},\mathcal{X})$
The theorem shows that $f_{imp}$ indeed generates imputations with minimum fairness (or maximum bias) based on Demographic Parity for a given classifier $f_{class}$.
@@
BFtS learns a classifier $f_{class}$ that minimizes the worst-case bias:
$\theta_{class}^{*}=\arg\min_{\theta_{class}}\sup_{\theta_{imp}}|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$
The optimal GNN classifier $f_{class}$, will achieve demographic parity ($\Delta DP=0$ in Sec 5) for the worst-case imputation (minimum fairness) $\hat{s}$ generated by $f_{imp}$. As $f_{class}$ is a minimax estimator, the maximal $\Delta DP$ of BFtS is minimum amongst all estimators of $s$.
@@
Let $s^{\prime}=f(\mathcal{G},\mathcal{X})$ be an imputation method independent of models $f_{class}$ and $f_{bias}$, then the BFtS imputation $\hat{s}=f_{imp}(\mathcal{G},\mathcal{X})$ is such that:
$JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\leq JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$
where $JS$ is the Jensen Shannon divergence.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is related to the convergence of adversarial learning. For independent imputation, if $s^{\prime}$ is inaccurate, then $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{bias}$ in Eq. 4 will be constant. The interplay between the three players makes BFtS more robust to convergence issues because the objective minimizes the upper bound on the JS divergence. This reduces the probability that the divergence vanishes during training (see details in the Appendix).
@@
5 Experimental Evaluation
We compare our approach (BFtS) against alternatives in terms of accuracy and fairness using real and synthetic data. We apply average precision (AVPR) and F1 for accuracy evaluation and $\Delta DP$ and $\Delta EQOP$ for fairness. Details about datasets, baselines, evaluation, and hyperparameters are provided in the Appendix. We group the baselines into two categories. First, we compare with methods that are explicitly designed to handle missing values: Debias Buyl & De Bie (2020), FairGNN Dai & Wang (2021), and RNF Du et al. (2021). Among these, RNF can operate when all sensitive attributes are missing, whereas Debias and FairGNN require access to at least some sensitive information. Both RNF and FairGNN rely on independent sensitive-attribute imputation procedures. Second, we compare our approach with state-of-the-art fair GNN models, including FairSIN Yang et al. (2024) and FairVGNN Wang et al. (2022). These methods cannot accommodate missing values directly, so we apply an independent imputation strategy consistent with the procedure used in FairGNN to supply the missing sensitive attributes before training. Additional experiments varying the GNN, scalability and complexity analysis using a synthetic large-scale graph, ablation studies (hyperparameter sensitivity, impact of LDAM loss, and impact of varying $\mathcal{V}_{S}$), and visualization of learnt representation and worst case assumption trade-off are also included in the Appendix. Our code can be found in an anonymous repository: https://anonymous.4open.science/r/BFtS-6ADA.
Figure 5: Fairness vs. accuracy results. The x-axis of each plot shows the F1 score. We plot $1-\Delta DP$ and $1-\Delta EQOP$ on the y-axis of rows 1 and 2, respectively. The top right corner of the plot represents a high F1 with low bias. BFtS often achieves better fairness for a similar value of F1.
@@ 5 Experimental Evaluation
Table 1: AVPR, F1, $\%\Delta DP$, and $\%\Delta EQOP$ without any sensitive information for BFtS and RNF (only baseline that operates in this setting). BFtS outperforms RNF in terms of fairness and accuracy.
@@ 5.1 Results and Analysis
Figure 4 shows the F1 score, $1-\Delta DP$, and $1-\Delta EQOP$ for different methods while varying their hyperparameters. The simulation graph was generated using a stochastic block model with different assortativity coefficients, i.e., the extent to which links exist within clusters compared with across clusters. Learning missing sensitive values and node labels under low assortativity is hard, and graphs with assortativity $0.17$ and $0.37$ represent the scenario described in Corollary 1, therefore, the JS divergence tends to be small for independent imputation, which may result in a lack of convergence for the baselines (FairGNN and Debias). As we increase the assortativity, all baselines can predict the labels, but BFtS still achieves a better fairness vs. accuracy trade-off in all cases.
We demonstrate the accuracy vs. fairness trade-off for real datasets in Figure 5. The top right corner of the plot represents high fairness and accuracy. Our model achieves better fairness and similar accuracy to the best baseline for the Bail, NBA, German, pokec-n and pokec-z. For all datasets, BFtS achieves a better fairness-accuracy tradeoff.
Table 1 shows the accuracy and fairness results for Bail, Credit, German, Pokec-z, Pokek-n and NBA without any sensitive attribute information. Among the baselines, only RNF works in this setting. BFtS outperforms RNF for the Bail, German, and Credit. BFtS also outperforms RNF in terms of fairness on Pokec-z,Pokec-n, and NBA with similar AVPR and F1.
BFtS outperforms the baselines because it effectively implements the worst-case assumption for missing value imputation. This assumption leads to stronger fairness guarantees (Theorem 3) than the baselines in cases where missing values are adversarial.
@@ 6 Conclusion
We investigate the challenge of incorporating fairness considerations into graph machine learning models when sensitive attributes are missing due to adversarial processes. Our solution is BFtS, a 3-player adversarial learning framework for the imputation of adversarially missing sensitive attributes that produce challenging values for graph-based fairness. Theoretical and empirical results demonstrate that BFtS achieves a better fairness $\times$ accuracy trade-off than existing alternatives.
Reproducibility statement
Details about datasets, baselines, evaluation, and hyperparameters are fully described in the Appendix. All real datasets and baselines applied in our experiments are publicly available. Our code can be found in an anonymous repository: https://anonymous.4open.science/r/BFtS-6ADA.
Acknowledgements
We acknowledge the support by the US Department of Transportation Tier-1 University Transportation Center (UTC) Transportation Cybersecurity Center for Advanced Research and Education (CYBER-CARE) (Grant No. 69A3552348332), and the Rice Ken Kennedy Institute.
@@ Appendix A Technical Appendices and Supplementary Material
Two of the models (or players) learned by our model are GNNs. Without loss of generality, we will assume that they are message-passing neural networks (Gilmer et al., 2017), which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
@@
Fairness Evaluation: We apply two bias metrics that will be defined based on a sensitive attribute $s$. Let $s=1$ be the sensitive group and $s=0$ be the non-sensitive group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute (Kamiran & Calders, 2009). We evaluate demographic parity as:
@@
Equality of Opportunity (EQOP) calculates the difference between true positive rates between sensitive and non-sensitive groups (Hardt et al., 2016) as follows:
@@
Equalized Odds (EQODDs) calculates the average difference between true positive rates and false positive rates between sensitive and non-sensitive groups (Grant, 2023) as follows:
$\begin{split}\Delta EQODDs(\hat{y},s,y)=\frac{1}{2}&|p(\hat{y}=1|s=1,y=1)-p(\hat{y}=1|s=0,y=1)|+\\ &|p(\hat{y}=1|s=1,y=0)-p(\hat{y}=1|s=0,y=0)|.\end{split}$ (10)
@@
We use $\Delta DP$ and $\Delta EQOP$ for most of our analysis as they are commonly used group-based metrics in fairness literature. Our theoretical analysis guarantees Demographic Parity, but we can change the $\mathcal{L}\_{bias}$ loss to account for a different type of fairness Madras et al. (2018).
@@
Bail (Jordan & Freiburger, 2015): Contains 18,876 nodes representing people who committed a criminal offense and were granted bail by US state courts between 1990 and 2009. Defendants are linked based on similarities in their demographics and criminal histories. The objective is to categorize defendants as likely to post bail vs. no bail. No-bail defendants are expected to be more likely to commit violent crimes if released. Race information is considered a sensitive attribute. Features include education, marital status, and profession.
Credit (Yeh & Lien, 2009): The dataset has $30,000$ nodes representing people each with $25$ features, including education, total overdue counts, and most recent payment amount. The objective is to forecast if a person will or won’t miss a credit card payment while considering age as a sensitive attribute. Nodes are connected based on payment patterns and spending similarities.
German (Hofmann, 1994): The dataset consists of 1000 nodes and 20 features (e.g., credit amount, savings) representing clients at a German bank. Labels are credit decisions for a person (good or bad credit). The sensitive attribute is gender. The goal is to predict who should receive credit.
NBA (Dai & Wang, 2021): Dataset based on a Kaggle dataset about the National Basketball Association (US). It contains 403 players as nodes connected based on their Twitter activity. Attributes represent player statistics and nationality is considered the sensitive attribute. The goal is to predict the players’ salaries (above/under the median).
Pokec (Takac & Zabovsky, 2012): This dataset has two variations: Pokec-z and pokec-n. The pokec dataset is based on a popular social network in Slovakia. The sensitive attribute in pokec-z and pokec-n are two different regions of Slovakia. The goal is to predict the working field of users.
@@
Debias (Zhang et al., 2018): Debias is an adversarial learning method that trains a classifier and an adversary simultaneously. The adversary is trained using the softmax output of the last layer of the classifier. It does not handle missing sensitive data. Hence, we only use $\mathcal{V}_{S}\in\mathcal{V}$ to calculate the adversarial loss.
FairGNN (Dai & Wang, 2021): FairGNN first uses a GNN to estimate the missing sensitive attributes. It then trains another GNN-based classifier and a DNN adversary. The adversary helps to eliminate bias from the representations learned by the GNN classifier.
RNF (Du et al., 2021): RNF imputes the missing sensitive attributes by first training the model with generalized cross entropy (Zhang & Sabuncu, 2018). It eliminates bias from the classification head. RNF uses samples with the same ground-truth label but distinct sensitive attributes and trains the classification head using their neutralized representations.
FairVGNN (Wang et al., 2022): FairVGNN learns fair node representations via automatically identifying and masking sensitive-correlated features. It requires complete sensitive information, so we use the same imputation method as FairGNN for the missing values.
FairSIN (Yang et al., 2024): FairSIN learns fair representations by adding additional fairness-facilitating features. It also requires complete sensitive information, so we use the same independent imputation method as FairGNN for the missing values. (Yang et al., 2024):
@@
The size of $\mathcal{V}_{S}$ was set to $30\%$ of $|\mathcal{V}|$ unless stated otherwise. Each GNN has two layers and a dropout probability of $0.5$. In practice, some sensitive information is often available, unless otherwise stated. Therefore, we concentrate the majority of our experiments on the assumption that there is a limited amount of sensitive information available, and we generate the imputations using the available sensitive information in Equation 2. The $\mathcal{V}_{S}$ was set to $30\%$ for all experiments unless stated otherwise. $\mathcal{V}_{L}$ is $300,400,1400,100,500,500$, and $800$ for the Simulation, German, Credit, NBA, pokec-z, pokec-n and Bail datasets, respectively. We adjust the values of $\alpha$ and $\beta$ for each dataset using cross-validation based on the F1 score. For all baselines, we choose the respective regularizers using cross-validation based on the same metrics. Every model, except RNF, combines the training of several networks with various convergence rates. We stop training RNF based on early stopping to eliminate overfitting. We train the rest of the models for certain epochs and choose the model with the best AVPR. All the results shown in the following section are the average of $10$ runs. We use GCN as the GNN architecture for all experiments unless mentioned otherwise.
@@
We will consider an adversarial active learning setting, where the goal of the adversary is to select $k$ labels (i.e., observed sensitive values) to minimize the bias $\mathcal{L}_{bias}$ of the data imputed by $f_{imp}$. More specifically, our proof will use the coverage model for active learning, where an unlabeled data point $x_{i}^{u}$ can be correctly predicted iff it is covered by at least one labeled point $x_{j}^{l}$. The goal is for the labeled points $\{x_{1}^{l},x_{2}^{l},\ldots x_{k}^{l}\}$ to cover as many unlabeled data points $\{x_{1}^{u},x_{2}^{u},\ldots x_{n-k}^{u}\}$ as possible (Yehuda et al., 2022; Ren et al., 2021). Thus, in our setting, the adversary’s goal is to select labels to minimize the coverage of the following sets of nodes: (1) nodes in the sensitive group ($s_{v}=1$) and the negative class ($y=0$) and (2) nodes in the non-sensitive group ($s_{v}=0$) and the positive class ($y=1$). This can be achieved by minimizing the coverage of such nodes. We provide a reduction from the minimum k-union problem.
Minimum k-union problem (MkU, Chlamtáč et al. (2017)): Given a set of sets $\mathcal{S}={S_{1},S_{2},\ldots S_{q}}$, where each set $S_{i}\subseteq\mathcal{I}$ and $\mathcal{I}$ is a set of items. The problem consists of selecting $r$ sets $S_{1},S_{2},\ldots S_{r}$ from $\mathcal{S}$ to minimize the coverage $S_{1}\cap S_{2}\cap\ldots S_{r}$.
Under our coverage model, the reduction is straightforward. Given an instance of MkU, we define a corresponding instance of the Adversarial Missingness against Data Bias (AMADB) problem as follows. We represent each item in $\mathcal{I}$ and each set in $\mathcal{S}$ as a data point $x$ and assign all data points to both the non-sensitive group ($s_{v}=0$) and the positive class ($y=1$). Moreover, we will add one data point $x^{\prime}$ that will belong to the sensitive group ($s_{v}=1$) and the positive class ($y=1$). Therefore, the resulting set of data points is $\{x_{1},x_{2},\ldots x_{n}\}$, where $n=|\mathcal{S}|+|\mathcal{I}|+1$. Moreover, let $x_{i}$ cover $x_{j}$ iff $x_{i}$ represents a set $S_{i}\in\mathcal{S}$, $x_{j}$ represents an item $i_{j}\in\mathcal{I}$, and $i_{j}\in S_{i}$. It follows that minimizing the bias in the AMABD instance by selecting data points $x_{j}$ to be labeled is equivalent to minimizing the coverage in the corresponding MkU instance.
@@
$\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}$ $[\log(f_{bias}(\mathbf{h}))]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{bias}(\mathbf{h}))]$
$=\sum_{h}p(\mathbf{h}|\hat{s}=1)\log(f_{bias}(\mathbf{h}))+\sum_{h}p(\mathbf{h}|\hat{s}=0)\log(1-f_{bias}(\mathbf{h}))$
If we fix $\theta_{class}$ and $\theta_{imp}$, then the optimal $f_{bias}$ is $\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}.$ By substituting the optimal $f_{bias}$ in Eq. 7 we get:
$\min_{\theta_{class}}\max_{\theta_{imp}}$~ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}]+$
@@
By removing the constants, we can further simplify the min-max optimization with optimal $f_{bias}$ to:
$\min_{\theta_{class}}\max_{\theta_{imp}}JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$
The optimal $f_{imp}$ thereby maximizes the following objective:
@@
Since $\hat{y}=\sigma(\mathbf{h.w})$, $f_{imp}$ generates imputation so that
$\min_{\theta_{class}}\sup_{\hat{s}}JS(p(\hat{y}|\hat{s}=1);p(\hat{y}|\hat{s}=0))$
The supremum value will make the two probabilities maximally different. Therefore, for the optimal $f_{imp}$ we get,
$\min_{\theta_{class}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Let $\Delta DP_{\theta_{class},\theta_{imp}}=|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$, then, the optimal $\theta_{imp}^{*}$ generates the worst-case (minimum fairness) imputation so that $\Delta DP_{\theta_{class},\theta_{imp}^{*}}\geq\Delta DP_{\theta_{class},\theta_{imp}}$